# RobustBlack: Challenging Black-Box Adversarial Attacks on State-of-the-Art Defenses

### v1

@@ Appendix C Increasing the Budget of Black-Box Attacks
Appendix D AutoAttack Robustness of LGV and GHOST Models
@@
To evaluate the impact of the LGV weight collection phase, and the GHOST perturbing phase on model robustness, we collected 10 models after finishing the first phase for both LGV and GHOST, prior to proceeding to the attacking phase. Specifically, the LGV variations were obtained by training the surrogate model for additional epochs using a high learning rate, during which the weights of the models were collected periodically. For GHOST networks, the variations were obtained by applying random perturbations to the skip connections of the surrogate model. Using the robust surrogate model WideResNet-101-2 introduced in (Peng et al., 2023) and described in Section 5, we evaluated the robustness of the resulting LGV models and compared it with the robustness of GHOST models. The robust accuracies for the collected models are reported in Table 9.
@@
Table 9: Robust accuracies of LGV and GHOST models before the attacking phase.
@@
Appendix E Detailed Tabular Results
We provide the tabular representation corresponding to the results presented in the main paper. Tables 10, 11, 12, 13 represent the success rates for section 3, 4, and 5 with vanilla surrogates and robust surrogates, respectively. These tables display the mean and standard deviation of success rates computed across three random seeds.
Table 10: Success rates of black-box attacks against one vanilla, and one robust ResNet50 model.
@@ Appendix E Detailed Tabular Results
Vanilla ResNet50 57.77 $\pm$0.00 71.31 $\pm$0.32 43.59 $\pm$ 0.17 66.79 $\pm$ 0.07 68.00 $\pm$ 0.18 99.95 $\pm$ 0.00 8.89 $\pm$ 0.20 64.53 $\pm$ 0.13 19.81 $\pm$ 1.46 81.08 $\pm$0.49 89.56 $\pm$0.09
Robust ResNet50 1.53 $\pm$ 0.00 1.94 $\pm$ 0.03 2.3 $\pm$ 0.01 1.66 $\pm$ 0.03 1.85 $\pm$ 0.04 2.54 $\pm$ 0.09 0.98 $\pm$ 0.10 1.68 $\pm$0.03 1.61 $\pm$0.10 2.83 $\pm$0.04 3.56 $\pm$0.06
Table 11: Success rates of white-box AutoAttack and black-box attacks against robust targets from Robustbench.
@@
21 Salman ResNet18 50.53 1.28 $\pm$0.00 1.41 $\pm$0.02 2.26 $\pm$ 0.04 1.24 $\pm$ 0.05 1.35 $\pm$ 0.02 1.68 $\pm$ 0.06 0.86 $\pm$ 0.10 1.25 $\pm$ 0.00 1.21 $\pm$0.18 1.63 $\pm$0.13 2.36 $\pm$0.04 4.06 $\pm$0.06
18 Salman ResNet50 43.84 1.68 $\pm$ 0.00 2.2 $\pm$ 0.10 2.81 $\pm$ 0.07 1.89 $\pm$ 0.04 1.72 $\pm$ 0.03 2.68 $\pm$ 0.13 1.02 $\pm$ 0.01 1.62 $\pm$ 0.00 1.69 $\pm$0.05 1.72 $\pm$0.26 2.83 $\pm$0.04 3.56 $\pm$0.06
17 Salman WideResNet50.2 42.87 1.34 $\pm$ 0.00 1.99 $\pm$ 0.04 2.78 $\pm$ 0.10 1.48 $\pm$ 0.05 1.59 $\pm$ 0.05 2.12 $\pm$ 0.08 0.93 $\pm$ 0.07 1.43 $\pm$ 0.00 1.49 $\pm$0.06 1.52 $\pm$0.08 2.35 $\pm$0.06 3.22 $\pm$0.06
12 Sing ViT-S, ConvStem 34.76 0.89 $\pm$ 0.00 1.41 $\pm$ 0.05 1.69 $\pm$ 0.05 0.97 $\pm$ 0.04 1.00 $\pm$ 0.00 1.51 $\pm$ 0.05 0.32 $\pm$ 0.06 0.95 $\pm$ 0.00 0.98 $\pm$0.03 0.72 $\pm$0.12 1.55 $\pm$0.00 1.54 $\pm$0.03
7 Liu ConvNeXt-B 28.01 0.63 $\pm$ 0.00 0.9 $\pm$ 0.02 0.97 $\pm$ 0.02 0.74 $\pm$ 0.02 0.74 $\pm$ 0.04 1.04 $\pm$ 0.04 0.19 $\pm$ 0.04 0.69 $\pm$ 0.00 0.63 $\pm$0.03 0.57 $\pm$0.04 1.38 $\pm$0.06 1.34 $\pm$0.04
5 Liu Swin-B 27.14 0.37 $\pm$ 0.00 0.63 $\pm$ 0.03 0.79 $\pm$ 0.07 0.41 $\pm$ 0.01 0.45 $\pm$ 0.02 0.72 $\pm$ 0.02 0.16 $\pm$ 0.02 0.37 $\pm$ 0.00 0.46 $\pm$0.07 0.36 $\pm$0.05 0.96 $\pm$0.02 1.14 $\pm$0.03
3 Liu ConvNeXt-L 25.95 0.49 $\pm$ 0.00 0.74 $\pm$ 0.09 1.07 $\pm$ 0.03 0.58 $\pm$ 0.03 0.49 $\pm$ 0.02 0.74 $\pm$ 0.01 0.18 $\pm$ 0.06 0.49 $\pm$ 0.00 0.49 $\pm$0.09 0.45 $\pm$0.05 1.16 $\pm$0.00 1.47 $\pm$0.01
2 Bai ConvNeXtV2-L, Swin-L 25.48 2.03 $\pm$ 0.00 2.65 $\pm$ 0.03 1.76 $\pm$ 0.05 2.40 $\pm$ 0.05 2.48 $\pm$ 0.03 2.55 $\pm$ 0.05 0.58 $\pm$ 0.05 2.00 $\pm$ 0.00 1.78 $\pm$0.16 1.01 $\pm$0.14 4.77 $\pm$0.13 2.94 $\pm$0.12
1 Liu Swin-L 25.98 0.56 $\pm$ 0.00 0.85 $\pm$ 0.04 1.14 $\pm$ 0.08 0.55 $\pm$ 0.01 0.64 $\pm$ 0.04 0.87 $\pm$ 0.00 0.21 $\pm$ 0.04 0.56 $\pm$ 0.00 0.49 $\pm$0.04 0.46 $\pm$0.04 1.26 $\pm$0.02 1.25 $\pm$0.01

### v2

@@ A-A Increasing the Computation Budget of Black-Box Attacks
Appendix B Increasing the Perturbation Budget of Black-Box Attacks
We have increased the perturbation budget to cover $\epsilon\in\{4/255,8/255,16/255\}$. In Table XIV, Table XV, and Table XVII, we only change the perturbation budget $\epsilon$.
Once we have increased the perturbation budget to $\epsilon=16/255$, we first study the impact of increasing the iteration budget in Table II, and Table XVI, then, we study the impact of increasing the query budget in Table III, and Table XVIII.
To further examine the impact of different defense strategies with increasing epsilon values and budgets, we present the results in the form of figures. Specifically, Figure 8 and Figure 9 correspond to epsilon values of 8/255 and 16/255, respectively. Additionally, Figure 10 for iteration and query budget of 2N and Figure 11 illustrate results for iteration budgets of 5N.
TABLE XIV: Attack Performance at Different Epsilon Values.
Attack Epsilon 4/255 Epsilon 8/255 Epsilon 16/255
MI-FGSM 57.77 $\rightarrow$ 1.53 77.33 $\rightarrow$ 3.53 88.79 $\rightarrow$ 10.52
DI-FGSM 71.31 $\rightarrow$ 1.94 87.07 $\rightarrow$ 4.18 94.55 $\rightarrow$ 12.11
TI-FGSM 43.59 $\rightarrow$ 2.30 66.35 $\rightarrow$ 4.68 83.53 $\rightarrow$ 15.61
VMI-FGSM 66.79 $\rightarrow$ 1.66 87.59 $\rightarrow$ 3.97 94.79 $\rightarrow$ 12.74
VNI-FGSM 68.00 $\rightarrow$ 1.85 90.10 $\rightarrow$ 4.09 96.94 $\rightarrow$ 13.21
ADMIX 77.90 $\rightarrow$ 2.54 93.04 $\rightarrow$ 4.68 98.28 $\rightarrow$ 14.24
UAP 8.89 $\rightarrow$ 0.98 25.70 $\rightarrow$ 2.37 70.78 $\rightarrow$ 9.93
GHOST 64.53 $\rightarrow$ 1.68 83.87 $\rightarrow$ 3.59 94.40 $\rightarrow$ 11.27
LGV 90.11 $\rightarrow$ 1.33 98.07 $\rightarrow$ 5.25 99.66 $\rightarrow$ 15.16
BASES 81.08 $\rightarrow$ 2.83 96.06 $\rightarrow$ 5.09 99.56 $\rightarrow$ 13.76
TREMBA 89.56 $\rightarrow$ 3.56 99.45 $\rightarrow$ 9.52 99.84 $\rightarrow$ 32.67
TABLE XV: Transfer Attack Performance Across Increased Epsilon 4/255, 8/255 and 16/255.
Model MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP SGM GHOST LGV
ResNet18 1.28 $\rightarrow$ 4.08 $\rightarrow$ 12.4 1.41 $\rightarrow$ 4.61 $\rightarrow$ 13.5 2.26 $\rightarrow$ 5.63 $\rightarrow$ 18.49 1.24 $\rightarrow$ 4.39 $\rightarrow$ 13.72 1.35 $\rightarrow$ 4.5 $\rightarrow$ 13.84 1.68 $\rightarrow$ 5.52 $\rightarrow$ 15.01 0.86 $\rightarrow$ 3.44 $\rightarrow$ 11.72 1.25 $\rightarrow$ 4.08 $\rightarrow$ 12.25 1.21 $\rightarrow$ 4.12 $\rightarrow$ 12.48 2.27 $\rightarrow$ 5.82 $\rightarrow$ 15.39
ResNet50 1.68 $\rightarrow$ 3.56 $\rightarrow$ 11.24 2.2 $\rightarrow$ 4.43 $\rightarrow$ 13.11 2.81 $\rightarrow$ 5.62 $\rightarrow$ 17.95 1.89 $\rightarrow$ 4.37 $\rightarrow$ 13.71 1.72 $\rightarrow$ 4.5 $\rightarrow$ 14.21 2.68 $\rightarrow$ 5.18 $\rightarrow$ 14.46 1.02 $\rightarrow$ 2.59 $\rightarrow$ 9.65 1.62 $\rightarrow$ 3.68 $\rightarrow$ 11.27 1.69 $\rightarrow$ 3.68 $\rightarrow$ 11.46 2.92 $\rightarrow$ 5.59 $\rightarrow$ 15.45
WideResNet50.2 1.34 $\rightarrow$ 3.82 $\rightarrow$ 10.02 1.99 $\rightarrow$ 4.6 $\rightarrow$ 12.09 2.78 $\rightarrow$ 6.06 $\rightarrow$ 17.27 1.48 $\rightarrow$ 4.52 $\rightarrow$ 12.5 1.59 $\rightarrow$ 4.66 $\rightarrow$ 13.02 2.12 $\rightarrow$ 5.13 $\rightarrow$ 12.94 0.93 $\rightarrow$ 2.62 $\rightarrow$ 7.95 1.43 $\rightarrow$ 3.79 $\rightarrow$ 9.96 1.49 $\rightarrow$ 3.67 $\rightarrow$ 10.52 2.58 $\rightarrow$ 5.39 $\rightarrow$ 13.92
ViT-S+ConvStem 0.89 $\rightarrow$ 2.36 $\rightarrow$ 6.26 1.41 $\rightarrow$ 2.98 $\rightarrow$ 7.76 1.69 $\rightarrow$ 3.34 $\rightarrow$ 10.38 0.97 $\rightarrow$ 2.73 $\rightarrow$ 8.37 1.0 $\rightarrow$ 2.75 $\rightarrow$ 8.51 1.51 $\rightarrow$ 3.09 $\rightarrow$ 8.43 0.32 $\rightarrow$ 1.31 $\rightarrow$ 4.56 0.95 $\rightarrow$ 2.28 $\rightarrow$ 6.37 0.98 $\rightarrow$ 2.31 $\rightarrow$ 6.59 1.53 $\rightarrow$ 3.11 $\rightarrow$ 8.93
ConvNeXt-B 0.63 $\rightarrow$ 1.56 $\rightarrow$ 4.2 0.9 $\rightarrow$ 1.85 $\rightarrow$ 4.88 0.97 $\rightarrow$ 2.35 $\rightarrow$ 7.5 0.74 $\rightarrow$ 1.85 $\rightarrow$ 5.81 0.74 $\rightarrow$ 1.82 $\rightarrow$ 5.83 1.04 $\rightarrow$ 2.51 $\rightarrow$ 5.6 0.19 $\rightarrow$ 0.74 $\rightarrow$ 2.53 0.69 $\rightarrow$ 1.5 $\rightarrow$ 4.3 0.63 $\rightarrow$ 1.48 $\rightarrow$ 4.22 0.92 $\rightarrow$ 2.48 $\rightarrow$ 6.1
Swin-B 0.37 $\rightarrow$ 1.35 $\rightarrow$ 3.72 0.63 $\rightarrow$ 1.59 $\rightarrow$ 4.3 0.79 $\rightarrow$ 2.15 $\rightarrow$ 6.93 0.41 $\rightarrow$ 1.62 $\rightarrow$ 5.18 0.45 $\rightarrow$ 1.7 $\rightarrow$ 5.2 0.72 $\rightarrow$ 2.2 $\rightarrow$ 5.55 0.16 $\rightarrow$ 0.72 $\rightarrow$ 2.81 0.37 $\rightarrow$ 1.33 $\rightarrow$ 3.74 0.46 $\rightarrow$ 1.25 $\rightarrow$ 3.95 0.91 $\rightarrow$ 1.65 $\rightarrow$ 5.31
ConvNeXt-L 0.49 $\rightarrow$ 1.5 $\rightarrow$ 4.35 0.74 $\rightarrow$ 2.07 $\rightarrow$ 5.18 1.07 $\rightarrow$ 2.57 $\rightarrow$ 7.39 0.58 $\rightarrow$ 1.97 $\rightarrow$ 6.45 0.49 $\rightarrow$ 2.07 $\rightarrow$ 6.56 0.74 $\rightarrow$ 2.41 $\rightarrow$ 6.25 0.18 $\rightarrow$ 0.78 $\rightarrow$ 2.31 0.49 $\rightarrow$ 1.61 $\rightarrow$ 4.28 0.49 $\rightarrow$ 1.37 $\rightarrow$ 4.64 1.07 $\rightarrow$ 2.44 $\rightarrow$ 5.93
ConvV2-L+Swin-L 2.03 $\rightarrow$ 3.46 $\rightarrow$ 6.23 2.65 $\rightarrow$ 4.75 $\rightarrow$ 7.67 1.76 $\rightarrow$ 3.51 $\rightarrow$ 7.52 2.4 $\rightarrow$ 4.38 $\rightarrow$ 9.37 2.48 $\rightarrow$ 4.85 $\rightarrow$ 9.52 2.55 $\rightarrow$ 4.58 $\rightarrow$ 8.83 0.58 $\rightarrow$ 1.51 $\rightarrow$ 3.44 2.0 $\rightarrow$ 3.61 $\rightarrow$ 6.38 1.78 $\rightarrow$ 3.69 $\rightarrow$ 6.9 2.57 $\rightarrow$ 4.8 $\rightarrow$ 9.15
Swin-L 0.56 $\rightarrow$ 1.25 $\rightarrow$ 3.43 0.85 $\rightarrow$ 1.84 $\rightarrow$ 4.46 1.14 $\rightarrow$ 2.43 $\rightarrow$ 6.84 0.55 $\rightarrow$ 1.51 $\rightarrow$ 5.07 0.64 $\rightarrow$ 1.74 $\rightarrow$ 5.28 0.87 $\rightarrow$ 2.02 $\rightarrow$ 5.12 0.21 $\rightarrow$ 0.59 $\rightarrow$ 2.48 0.56 $\rightarrow$ 1.31 $\rightarrow$ 3.41 0.49 $\rightarrow$ 1.18 $\rightarrow$ 3.61 1.03 $\rightarrow$ 1.92 $\rightarrow$ 4.84
TABLE XVI: Transfer Attack Performance Across Different Iteration Counts N, 2N, 5N For Epsilon 16/255.
Model MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP SGM GHOST LGV
ResNet18 12.4 $\rightarrow$ 13.38 $\rightarrow$ 14.63 13.5 $\rightarrow$ 15.12 $\rightarrow$ 15.77 18.49 $\rightarrow$ 20.23 $\rightarrow$ 20.76 13.72 $\rightarrow$ 15.65 $\rightarrow$ 16.33 13.84 $\rightarrow$ 15.77 $\rightarrow$ 16.22 15.01 $\rightarrow$ 15.24 $\rightarrow$ 16.22 11.72 $\rightarrow$ 11.72 $\rightarrow$ 11.72 12.25 $\rightarrow$ 13.57 $\rightarrow$ 14.37 12.48 $\rightarrow$ 14.14 $\rightarrow$ 15.16 15.39 $\rightarrow$ 17.05 $\rightarrow$ 17.47
ResNet50 11.24 $\rightarrow$ 12.11 $\rightarrow$ 13.27 13.11 $\rightarrow$ 13.92 $\rightarrow$ 14.89 17.95 $\rightarrow$ 19.29 $\rightarrow$ 20.32 13.71 $\rightarrow$ 14.99 $\rightarrow$ 15.24 14.21 $\rightarrow$ 15.39 $\rightarrow$ 15.52 14.46 $\rightarrow$ 15.08 $\rightarrow$ 15.24 9.65 $\rightarrow$ 9.65 $\rightarrow$ 9.65 11.27 $\rightarrow$ 12.46 $\rightarrow$ 13.18 11.46 $\rightarrow$ 12.86 $\rightarrow$ 14.02 15.45 $\rightarrow$ 17.2 $\rightarrow$ 16.98
WideResNet50.2 10.02 $\rightarrow$ 10.43 $\rightarrow$ 11.07 12.09 $\rightarrow$ 12.7 $\rightarrow$ 13.08 17.27 $\rightarrow$ 18.88 $\rightarrow$ 19.11 12.5 $\rightarrow$ 13.46 $\rightarrow$ 12.88 13.02 $\rightarrow$ 13.87 $\rightarrow$ 13.55 12.94 $\rightarrow$ 12.38 $\rightarrow$ 12.5 7.95 $\rightarrow$ 7.95 $\rightarrow$ 7.95 9.96 $\rightarrow$ 10.66 $\rightarrow$ 11.07 10.52 $\rightarrow$ 11.1 $\rightarrow$ 11.56 13.92 $\rightarrow$ 14.24 $\rightarrow$ 14.39
ViT-S+ConvStem 6.26 $\rightarrow$ 6.59 $\rightarrow$ 6.73 7.76 $\rightarrow$ 8.32 $\rightarrow$ 8.15 10.38 $\rightarrow$ 10.38 $\rightarrow$ 10.54 8.37 $\rightarrow$ 9.01 $\rightarrow$ 8.46 8.51 $\rightarrow$ 8.93 $\rightarrow$ 8.6 8.43 $\rightarrow$ 8.43 $\rightarrow$ 8.09 4.56 $\rightarrow$ 4.56 $\rightarrow$ 4.56 6.37 $\rightarrow$ 6.68 $\rightarrow$ 6.87 6.59 $\rightarrow$ 6.82 $\rightarrow$ 7.04 8.93 $\rightarrow$ 9.18 $\rightarrow$ 8.54
ConvNeXt-B 4.2 $\rightarrow$ 4.12 $\rightarrow$ 4.3 4.88 $\rightarrow$ 5.17 $\rightarrow$ 5.12 7.5 $\rightarrow$ 8.47 $\rightarrow$ 8.13 5.81 $\rightarrow$ 6.04 $\rightarrow$ 5.54 5.83 $\rightarrow$ 5.94 $\rightarrow$ 5.49 5.6 $\rightarrow$ 5.3 $\rightarrow$ 5.28 2.53 $\rightarrow$ 2.53 $\rightarrow$ 2.53 4.3 $\rightarrow$ 4.28 $\rightarrow$ 4.22 4.22 $\rightarrow$ 4.35 $\rightarrow$ 4.57 6.1 $\rightarrow$ 6.04 $\rightarrow$ 6.12
Swin-B 3.72 $\rightarrow$ 3.9 $\rightarrow$ 4.06 4.3 $\rightarrow$ 4.43 $\rightarrow$ 5.18 6.93 $\rightarrow$ 7.3 $\rightarrow$ 7.4 5.18 $\rightarrow$ 5.36 $\rightarrow$ 5.2 5.2 $\rightarrow$ 5.57 $\rightarrow$ 5.33 5.55 $\rightarrow$ 5.12 $\rightarrow$ 4.94 2.81 $\rightarrow$ 2.81 $\rightarrow$ 2.81 3.74 $\rightarrow$ 3.95 $\rightarrow$ 4.03 3.95 $\rightarrow$ 4.01 $\rightarrow$ 4.22 5.31 $\rightarrow$ 5.6 $\rightarrow$ 5.65
ConvNeXt-L 4.35 $\rightarrow$ 4.02 $\rightarrow$ 4.09 5.18 $\rightarrow$ 5.13 $\rightarrow$ 5.52 7.39 $\rightarrow$ 7.62 $\rightarrow$ 7.98 6.45 $\rightarrow$ 6.14 $\rightarrow$ 5.52 6.56 $\rightarrow$ 6.4 $\rightarrow$ 5.7 6.25 $\rightarrow$ 5.62 $\rightarrow$ 5.6 2.31 $\rightarrow$ 2.31 $\rightarrow$ 2.31 4.28 $\rightarrow$ 4.02 $\rightarrow$ 4.15 4.64 $\rightarrow$ 4.2 $\rightarrow$ 4.41 5.93 $\rightarrow$ 6.12 $\rightarrow$ 5.62
ConvV2-L+Swin-L 6.23 $\rightarrow$ 6.41 $\rightarrow$ 6.51 7.67 $\rightarrow$ 8.31 $\rightarrow$ 8.43 7.52 $\rightarrow$ 7.99 $\rightarrow$ 8.06 9.37 $\rightarrow$ 9.2 $\rightarrow$ 8.61 9.52 $\rightarrow$ 9.55 $\rightarrow$ 8.98 8.83 $\rightarrow$ 8.34 $\rightarrow$ 8.19 3.44 $\rightarrow$ 3.44 $\rightarrow$ 3.44 6.38 $\rightarrow$ 6.36 $\rightarrow$ 6.6 6.9 $\rightarrow$ 6.73 $\rightarrow$ 6.73 9.15 $\rightarrow$ 9.03 $\rightarrow$ 8.9
Swin-L 3.43 $\rightarrow$ 3.53 $\rightarrow$ 3.61 4.46 $\rightarrow$ 4.38 $\rightarrow$ 4.51 6.84 $\rightarrow$ 7.32 $\rightarrow$ 6.97 5.07 $\rightarrow$ 4.94 $\rightarrow$ 4.64 5.28 $\rightarrow$ 5.25 $\rightarrow$ 4.92 5.12 $\rightarrow$ 4.53 $\rightarrow$ 4.43 2.48 $\rightarrow$ 2.48 $\rightarrow$ 2.48 3.41 $\rightarrow$ 3.51 $\rightarrow$ 3.53 3.61 $\rightarrow$ 3.51 $\rightarrow$ 3.53 4.84 $\rightarrow$ 4.89 $\rightarrow$ 4.92
TABLE XVII: Query Attack Performance Across Increased Epsilon 4/255, 8/255 and 16/255.
Model BASES TREMBA
ResNet18 2.36 $\rightarrow$ 5.33 $\rightarrow$ 13.43 4.06 $\rightarrow$ 11.3 $\rightarrow$ 37.39
ResNet50 2.83 $\rightarrow$ 5.09 $\rightarrow$ 13.76 3.56 $\rightarrow$ 9.52 $\rightarrow$ 32.67
WideResNet50.2 2.35 $\rightarrow$ 5.27 $\rightarrow$ 12.07 3.22 $\rightarrow$ 9.04 $\rightarrow$ 31.77
ViT-S+ConvStem 1.55 $\rightarrow$ 3.39 $\rightarrow$ 7.93 1.54 $\rightarrow$ 5.42 $\rightarrow$ 24.11
ConvNeXt-B 1.38 $\rightarrow$ 2.53 $\rightarrow$ 5.8 1.34 $\rightarrow$ 4.46 $\rightarrow$ 18.54
Swin-B 0.96 $\rightarrow$ 2.26 $\rightarrow$ 5.2 1.14 $\rightarrow$ 4.11 $\rightarrow$ 17.17
ConvNeXt-L 1.16 $\rightarrow$ 2.7 $\rightarrow$ 6.04 1.47 $\rightarrow$ 4.43 $\rightarrow$ 17.21
ConvV2-L+Swin-L 4.77 $\rightarrow$ 6.73 $\rightarrow$ 10.12 2.94 $\rightarrow$ 7.59 $\rightarrow$ 21.84
Swin-L 1.26 $\rightarrow$ 2.18 $\rightarrow$ 4.79 1.25 $\rightarrow$ 4.02 $\rightarrow$ 16.41
TABLE XVIII: Query Attack Performance Across Different Query Counts N, 2N For Epsilon 16/255.
Model BASES TREMBA
ResNet18 13.43 $\rightarrow$ 13.73 37.39 $\rightarrow$ 43.06
ResNet50 13.76 $\rightarrow$ 14.05 32.67 $\rightarrow$ 37.78
WideResNet50.2 12.07 $\rightarrow$ 12.23 31.77 $\rightarrow$ 36.24
ViT-S+ConvStem 7.93 $\rightarrow$ 8.07 24.11 $\rightarrow$ 29.45
ConvNeXt-B 5.8 $\rightarrow$ 5.93 18.54 $\rightarrow$ 22.51
Swin-B 5.2 $\rightarrow$ 5.2 17.17 $\rightarrow$ 21.44
ConvNeXt-L 6.04 $\rightarrow$ 6.17 17.21 $\rightarrow$ 21.02
ConvV2-L+Swin-L 10.12 $\rightarrow$ 10.29 21.84 $\rightarrow$ 26.34
Swin-L 4.79 $\rightarrow$ 4.99 16.41 $\rightarrow$ 20.46
Figure 8: Success rate of blackbox attacks against SoTA defenses for epsillon 8/255.
Figure 9: Success rate of blackbox attacks against SoTA defenses for epsilon 16/255.
Figure 10: Success rate of blackbox attacks against SoTA defenses for epsilon 16/255 and iteration=2N.
Figure 11: Success rate of blackbox attacks against SoTA defenses for epsilon 16/255 and iteration=5N.
B-A AutoAttack Robustness of LGV and GHOST Models
@@
To evaluate the impact of the LGV weight collection phase, and the GHOST perturbing phase on model robustness, we collected 10 models after finishing the first phase for both LGV and GHOST, prior to proceeding to the attacking phase. Specifically, the LGV variations were obtained by training the surrogate model for additional epochs using a high learning rate, during which the weights of the models were collected periodically. For GHOST networks, the variations were obtained by introducing a random perturbation to the scaling factor in the skip connections. Using the robust surrogate model WideResNet-101-2 introduced in [54] and described in Section V, we evaluated the robustness of the resulting LGV models and compared it with the robustness of GHOST models. The robust accuracies for the collected models are reported in Table XIX.
@@
TABLE XIX: Robust accuracies of LGV and GHOST models before the attacking phase.
@@
B-B Detailed Tabular Results
We provide the tabular representation corresponding to the results presented in the main paper. Tables XX, XXI, XXII, XXIII represent the success rates for section III, IV, and V with vanilla surrogates and robust surrogates, respectively. These tables display the mean and standard deviation of success rates computed across three random seeds.
B-C Reproduction Package and Availability
The source code to reproduce the experiments of this paper is available at https://github.com/serval-uni-lu/RobustBlack.git.
TABLE $X$: Success rates of black-box attacks against one vanilla, and one robust ResNet50 model.
@@ B-C Reproduction Package and Availability
Vanilla ResNet50 57.77 $\pm$0.00 71.31 $\pm$0.32 43.59 $\pm$ 0.17 66.79 $\pm$ 0.07 68.00 $\pm$ 0.18 99.95 $\pm$ 0.00 8.89 $\pm$ 0.20 64.53 $\pm$ 0.13 90.11 $\pm$ 0.12 81.08 $\pm$0.49 89.56 $\pm$0.09
Robust ResNet50 1.53 $\pm$ 0.00 1.94 $\pm$ 0.03 2.3 $\pm$ 0.01 1.66 $\pm$ 0.03 1.85 $\pm$ 0.04 2.54 $\pm$ 0.09 0.98 $\pm$ 0.10 1.68 $\pm$0.03 1.33 $\pm$0.01 2.83 $\pm$0.04 3.56 $\pm$0.06
TABLE XXI: Success rates of white-box AutoAttack and black-box attacks against robust targets from Robustbench.
@@
21 Salman ResNet18 50.53 1.28 $\pm$0.00 1.41 $\pm$0.02 2.26 $\pm$ 0.04 1.24 $\pm$ 0.05 1.35 $\pm$ 0.02 1.68 $\pm$ 0.06 0.86 $\pm$ 0.10 1.25 $\pm$ 0.00 1.21 $\pm$0.18 2.27 $\pm$0.11 2.36 $\pm$0.04 4.06 $\pm$0.06
18 Salman ResNet50 43.84 1.68 $\pm$ 0.00 2.2 $\pm$ 0.10 2.81 $\pm$ 0.07 1.89 $\pm$ 0.04 1.72 $\pm$ 0.03 2.68 $\pm$ 0.13 1.02 $\pm$ 0.01 1.62 $\pm$ 0.00 1.69 $\pm$0.05 2.92 $\pm$0.01 2.83 $\pm$0.04 3.56 $\pm$0.06
17 Salman WideResNet50.2 42.87 1.34 $\pm$ 0.00 1.99 $\pm$ 0.04 2.78 $\pm$ 0.10 1.48 $\pm$ 0.05 1.59 $\pm$ 0.05 2.12 $\pm$ 0.08 0.93 $\pm$ 0.07 1.43 $\pm$ 0.00 1.49 $\pm$0.06 2.58 $\pm$0.09 2.35 $\pm$0.06 3.22 $\pm$0.06
12 Sing ViT-S, ConvStem 34.76 0.89 $\pm$ 0.00 1.41 $\pm$ 0.05 1.69 $\pm$ 0.05 0.97 $\pm$ 0.04 1.00 $\pm$ 0.00 1.51 $\pm$ 0.05 0.32 $\pm$ 0.06 0.95 $\pm$ 0.00 0.98 $\pm$0.03 1.53 $\pm$0.08 1.55 $\pm$0.00 1.54 $\pm$0.03
7 Liu ConvNeXt-B 28.01 0.63 $\pm$ 0.00 0.9 $\pm$ 0.02 0.97 $\pm$ 0.02 0.74 $\pm$ 0.02 0.74 $\pm$ 0.04 1.04 $\pm$ 0.04 0.19 $\pm$ 0.04 0.69 $\pm$ 0.00 0.63 $\pm$0.03 0.92 $\pm$0.07 1.38 $\pm$0.06 1.34 $\pm$0.04
5 Liu Swin-B 27.14 0.37 $\pm$ 0.00 0.63 $\pm$ 0.03 0.79 $\pm$ 0.07 0.41 $\pm$ 0.01 0.45 $\pm$ 0.02 0.72 $\pm$ 0.02 0.16 $\pm$ 0.02 0.37 $\pm$ 0.00 0.46 $\pm$0.07 0.91 $\pm$0.03 0.96 $\pm$0.02 1.14 $\pm$0.03
3 Liu ConvNeXt-L 25.95 0.49 $\pm$ 0.00 0.74 $\pm$ 0.09 1.07 $\pm$ 0.03 0.58 $\pm$ 0.03 0.49 $\pm$ 0.02 0.74 $\pm$ 0.01 0.18 $\pm$ 0.06 0.49 $\pm$ 0.00 0.49 $\pm$0.09 1.07 $\pm$0.03 1.16 $\pm$0.00 1.47 $\pm$0.01
2 Bai ConvNeXtV2-L, Swin-L 25.48 2.03 $\pm$ 0.00 2.65 $\pm$ 0.03 1.76 $\pm$ 0.05 2.40 $\pm$ 0.05 2.48 $\pm$ 0.03 2.55 $\pm$ 0.05 0.58 $\pm$ 0.05 2.00 $\pm$ 0.00 1.78 $\pm$0.16 2.57 $\pm$0.10 4.77 $\pm$0.13 2.94 $\pm$0.12
1 Liu Swin-L 25.98 0.56 $\pm$ 0.00 0.85 $\pm$ 0.04 1.14 $\pm$ 0.08 0.55 $\pm$ 0.01 0.64 $\pm$ 0.04 0.87 $\pm$ 0.00 0.21 $\pm$ 0.04 0.56 $\pm$ 0.00 0.49 $\pm$0.04 1.03 $\pm$0.03 1.26 $\pm$0.02 1.25 $\pm$0.01